ALL RIGHTS RESERVED.
CEBS - GBA 1 - Practice Exam Questions
And Answers 100% Verified.
Which of the following statements correctly describes minimum requirements that group benefit
plans typically must meet in order to be prospectively rated on contract renewal? -
Answers✔The more unpredictable the event that will trigger a claim, the larger the minimum
group size and premium requirements
Which of the following statements correctly describes the term graduation as it applies to
mortality tables? - Answers✔It is a mathematical process that removes anomalies that should not
appear in the mortality data
If the plan sponsor pays the premium, which of the following group insured benefits is taxable to
the plan member? - Answers✔Long-term disability (LTD)
Which of the following statements correctly describes phases of a group plan marketing? -
Answers✔Plan features of both contributory and noncontributory benefit plans must be
communicated to plan members during the enrollment process
Which of the following statements correctly describes a characteristic of an insurable risk? -
Answers✔The loss must have financial significance to the insured (The loss must not be
catastrophic to the insurer. Option C: The loss must occur by chance. Option D: The loss must be
predictable for the insurer.)
With respect to marketing a group insurance plan, which of the following statements best
describes insurer representatives known as account executives? - Answers✔They are involved in
enrolling group members and installing new plans
Which of the following statements correctly describes the statistical concept known as the law of
large numbers? - Answers✔Future probabilities of loss are not valid unless a large number of
incidents exist to ensure the outcomes are less likely to deviate
Which of the following statements correctly describes an insured plan that is insurer-
administered? - Answers✔For self-insured plans with an administrative services only (ASO)
arrangement, the insurer bills the plan sponsor monthly for administration fees
, ©THESTAR EXAM SOLUTIONS 2024/2025
ALL RIGHTS RESERVED.
If the probability that an individual will live one year is 0.95 and the probability of dying in the
following year is 0.10, what is the probability that an individual will survive both years? -
Answers✔0.855 (P (Survive both years) = (0.95) x (1 - 0.1) = 0.95 x 0.90 = 0.855.)
Which of the following statements correctly describes the group benefits enrollment process for
new employees? - Answers✔An employee can enroll in the plan any time between his or her
date of hire and the end of a waiting period
A type of service offered under a third-party administrators (TPAs) claims settlement only
arrangement is: - Answers✔Adjudicating and paying claims
Unless otherwise specified in the group insurance contract, a plan member's status under a group
benefits plan terminates the earliest at which: - Answers✔The plan member retires (Unless
otherwise specified in the group insurance contract, the order in which coverage terminates is
effective the earliest date the plan member retires, the plan member's service with the plan
sponsor is terminated, the plan member ceases active work, and the plan member is no longer a
member of an eligible class and premium payments stop.)
Which of the following statements correctly describes eligibility requirements under most group
insurance plans? - Answers✔Unmarried children under a specified age and dependent on the
plan member for support are eligible for coverage
Which of the following statements correctly describes the negotiated placement method of
marketing a group benefits plan? - Answers✔Once potential insurers are identified, each
company submits a brief report that highlights its general strategy, philosophy and proposed
group benefits plan
True North Insurance has established a waiver-of-premium reserve of $100,000 and will have to
pay out an optional life insurance benefit of $200,000. With an interest rate of 6% compounded
annually, how many years will it take for the reserve of $100,000 to accumulate to $200,000? -
Answers✔12 years (Use the relationship between future value and present value to solve for the
number of time periods.
FVt = PV x (1 + r)^t
$200,000 = $100,000 x (1 + 6%)^t
$200,000/$100,000 = (1.06)^t
2.00 = (1.06)^t
The number of time periods can be approximated using the future value table. With an interest
rate of 6%, it will take 12 years to accumulate to a sum of $200,000 (or 2.01220).)
Which of the following statements correctly describes a group insurance policy? - Answers✔It is
issued on a group of individuals for protection against financial hazards, such as death, accident
and sickness